# Theory and Design of Space-Time Metallic Metasurfaces for Wireless Communications

### v1

@@ III-B Diffracted Waves
III-C Frequency Mixing
Figure 8: FDTD results for a C-G frequency mixer for an incident wave of frequency $\omega_{0}=30$ GHz: (a) $T_{\text{M}}=T_{0}$; (b) $T^{\prime}_{\text{M}}=2T_{0}$; (c) $T^{\prime}_{\text{M}}=3T_{0}$. Normal TM incidence is considered. Parameters: $t_{C}=[0,\ 0.25T_{\text{M}}]$, $t_{G}=[0.25T_{\text{M}},\ T_{\text{M}}]$, $P=1.2\lambda_{0}$, $W=0.5P$.
@@ III-D Transmissive Beamformer
In contrast, Figs. 9(c)-(d) depict the phenomenology when the temporal period is modified. For simplicity, TE normal incidence is assumed and slit width is increased to $W=0.6P$ in order to achieve a higher transmission. Furthermore, a spatial period greater than the wavelength of the incident wave is imposed to enrich the diffraction diagram since the first (spatial) grating lobes are excited. Fig. 9(c) shows the diffraction pattern for the same time period as considered throughout the previous section of the manuscript ($T_{\text{M}}=4T_{0}$). In this case, a symmetric pattern is appreciated where the temporal $(0,n)$-th harmonics are in the same direction ($\theta_{0n}=0^{\mathrm{o}}$). Moreover, all the $(0,n)$-th harmonics are propagative due to their temporal nature and the normal incidence conditions. However, since spatial modulation has been included, new spatiotemporal harmonics emerge symmetrically at both the upper and lower parts of the diagram. Finally, in Fig. 9(d), it is observed how the higher-order $(m,n)$-th harmonics approach to the normal when the temporal period is reduced (from $T_{\text{M}}=4T_{0}$ to $T_{\text{M}}=T_{0}$), as it was predicted in Section II-D.
Table I provides a comparison of the diffraction angles for the previous cases shown in Fig. 9. They have been extracted with the analytical approach [by means of eq. (15b)] and the FDTD method. A good agreement is appreciated between the analytical and numerical results.
TABLE I: Diffraction angles obtained for each $(m,n)$-th harmonic using the FDTD and analytical method for the cases shown in Fig. 9.
@@
Frequency-mixing phenomena can be appreciated in Figs. 9(c) and 9(d) as well. As discussed in Section III-C, two different space-time modulations can lead to different output frequencies while maintaining the same diffraction angle. This is case for the diffraction orders ($m=1$, $n=1$) in Fig. 9(c) and ($m^{\prime}=4$, $n^{\prime}=4$) in Fig. 9(d). Both diffraction orders have associated the same diffraction angle ($\theta_{11}=\theta_{4^{\prime}4^{\prime}}=41.77^{\mathrm{o}}$) but different output frequencies ($\omega_{n=1}=1.25\omega_{0}$, $\omega_{n^{\prime}=4}=5\omega_{0}$). In a general case, eqs. (37) and (38) can be applied in the design of frequency mixers under oblique and normal incidence, respectively, by simply tuning the space-time parameters of the modulation. The key parameter in design is the time period, as the spatial period cannot be easily reconfigured electronically in the current metastructure. The present discussion opens the way to the design of spatiotemporal frequency mixers based on aperiodic metallic structures, which are expected to provide more degrees of freedom in design and ease of control than purely-periodic metallic ones [32].
III-E Reflective Beamformer
@@
Fig. 10(a) shows the curves obtained by the proposed analytical method for the reflection coefficient $|S_{11}|$ when varying the duty cycle $D$. In this case, $DT_{\text{M}}$ represents the time in which the C-G metasurface is in the conductor state. FDTD simulations of the electric-field distributions are also included as insets. The black, blue and red solid lines are associated to the duty cycle $D=0.75$, $D=0.5$ and $D=0.25$, respectively. It can be seen how the amplitude of the reflection coefficient decreases as the space-time metasurface remains less time in the conductor (fully reflective) state; namely, as $D$ decreases. This is in line with the former theoretical discussion and with the FDTD results inserted in Fig. 10(a). Moreover, the oblique incident angle ($\theta_{0}=20^{\mathrm{o}}$) leads to the emerging of the grating lobe approximately at 23 GHz. This phenomenology is well caught by the circuital approach. It can be also appreciated in the FDTD results, as the grating lobes make appearance at 30 GHz (see the lower half space) but do not show at 20 GHz, below the grating-lobe regime. Finally, Fig. 10(b) shows the amplitude of the reflection coefficient in the case of varying the spatial ratio $W/P$ and fixing the duty cycle. In this case, less reflection is observed as the spatial ratio (slit width $W$) increases. As Figs. 10(a) and (10)(b) illustrate,
@@
A crucial difference related to these two parameters should be noted. The modification of $D$ and $W$ is useful to adjust the reflection coefficient of the system, as explained in this section. However, as discussed in Section III-A, the harmonic distribution of the diffraction pattern is also affected. Nevertheless, the majority of $(m,n)$-th harmonics that can attain significant amplitudes when varying $W$ are of evanescent nature. This is not an issue from the beamforming perspective, since this parameter remains fixed in the grating state. For this reason, this parameter is the best choice for setting the desired reflection coefficient of the system. On the other hand, when $D$ is modified, many of these $(m,n)$th harmonics that can attain significant amplitudes will be propagative. Thus, this property of the temporal parameters turns them into the best option for adjusting the beamsteering of the system.
@@ Acknowledgments
This work has been supported by grant PID2020-112545RB-C54 funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. It has also been supported by grants PDC2022-133900-I00, TED2021-129938B-I00 and TED2021-131699B-I00, and by Ministerio de Universidades and the European Union NextGenerationEU, under Programa Margarita Salas, by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR, and by IJC2020-043599-I funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEUPRTR.

### v2

@@ III.2 Diffracted Waves
III.3 Frequency Mixing
Figure 8: FDTD results for a C-G frequency mixer for an incident wave of frequency $\omega_{0}$ = $2\pi\cdot 30$ GHz: (a) $T_{\text{M}}=T_{0}$; (b) $T^{\prime}_{\text{M}}=2T_{0}$; (c) $T^{\prime}_{\text{M}}=3T_{0}$. Normal TM incidence is considered. Parameters: $t_{C}=[0,\ 0.25T_{\text{M}}]$, $t_{G}=[0.25T_{\text{M}},\ T_{\text{M}}]$, $P=1.2\lambda_{0}$, $W=0.5P$.
@@ III.4 Transmissive Beamformer
In contrast, Figs. 9(c)-(d) depict the phenomenology when the temporal period is modified. For simplicity, TE normal incidence is assumed and slit width is increased to $W=0.6P$ in order to achieve a higher transmission. Furthermore, a spatial period greater than the wavelength of the incident wave is imposed to enrich the diffraction diagram since the first (spatial) grating lobes are excited. Fig. 9(c) shows the diffraction pattern for the same time period as considered throughout the previous section of the manuscript ($T_{\text{M}}=4T_{0}$). In this case, a symmetric pattern is appreciated where the temporal $(0,n)$-th harmonics are in the same direction ($\theta_{0n}=0^{\mathrm{o}}$). Moreover, all the $(0,n)$-th harmonics are propagative due to their temporal nature and the normal incidence conditions. However, since spatial modulation has been included, new spatiotemporal harmonics emerge symmetrically at both the upper and lower parts of the diagram. Finally, in Fig. 9(d), it is observed how the higher-order $(m,n)$-th harmonics approach to the normal when the temporal period is reduced (from $T_{\text{M}}=4T_{0}$ to $T_{\text{M}}=T_{0}$), as it was predicted in Section II.4.
Table 1 provides a comparison of the diffraction angles for the previous cases shown in Fig. 9. They have been extracted with the analytical approach [by means of eq. (15b)] and the FDTD method. A good agreement is appreciated between the analytical and numerical results.
Table 1: Diffraction angles obtained for each $(m,n)$-th harmonic using the FDTD and analytical method for the cases shown in Fig. 9.
@@
Frequency-mixing phenomena can be appreciated in Figs. 9(c) and 9(d) as well. As discussed in Section III.3, two different space-time modulations can lead to different output frequencies while maintaining the same diffraction angle. This is case for the diffraction orders ($m=1$, $n=1$) in Fig. 9(c) and ($m^{\prime}=4$, $n^{\prime}=4$) in Fig. 9(d). Both diffraction orders have associated the same diffraction angle ($\theta_{11}=\theta_{4^{\prime}4^{\prime}}=41.77^{\mathrm{o}}$) but different output frequencies ($\omega_{n=1}=1.25\omega_{0}$, $\omega_{n^{\prime}=4}=5\omega_{0}$). In a general case, eqs. (37) and (38) can be applied in the design of frequency mixers under oblique and normal incidence, respectively, by simply tuning the space-time parameters of the modulation. The key parameter in design is the time period, as the spatial period cannot be easily reconfigured electronically in the current metastructure. Without loss of generality, with this scenario, achieving the demodulation of the same signal by users located in different positions would be possible, which is highly interesting for applications in mobile or satellite communications. The present discussion opens the way to the design of spatiotemporal frequency mixers based on aperiodic metallic structures, which are expected to provide more degrees of freedom in design and ease of control than purely-periodic metallic ones [33].
III.5 Reflective Beamformer
@@
Fig. 10(a) shows the curves obtained by the proposed analytical method for the reflection coefficient $|S_{11}|$ when varying the duty cycle $D$. In this case, $DT_{\text{M}}$ represents the time in which the C-G metasurface is in the conductor state. FDTD simulations of the electric-field distributions are also included as insets. The black, blue and red solid lines are associated to the duty cycle $D=0.75$, $D=0.5$ and $D=0.25$, respectively. It can be seen how the amplitude of the reflection coefficient decreases as the space-time metasurface remains less time in the conductor (fully reflective) state; namely, as $D$ decreases. This is in line with the former theoretical discussion and with the FDTD results inserted in Fig. 10(a). Moreover, the oblique incident angle ($\theta_{0}=20^{\mathrm{o}}$) leads to the emerging of the grating lobe approximately at 23 GHz. This phenomenology is well caught by the circuital approach. It can be also appreciated in the FDTD results, as the grating lobes make appearance at 30 GHz (see the lower half space) but do not show at 20 GHz, below the grating-lobe regime. Finally, Fig. 10(b) shows the amplitude of the reflection coefficient in the case of varying the spatial ratio $W/P$ and fixing the duty cycle. In this case, less reflection is observed as the spatial ratio (slit width $W$) increases.
@@
A crucial difference related to these two parameters should be noted. The modification of $D$ and $W$ is useful to adjust the reflection coefficient of the system, as explained in this section. However, as discussed in Section III.1, the harmonic distribution of the diffraction pattern is also affected. Nevertheless, the majority of $(m,n)$-th harmonics that can attain significant amplitudes when varying $W$ are of evanescent nature. This is not an issue from the beamforming perspective, since this parameter remains fixed in the grating state. On the other hand, when $D$ is modified, many of these $(m,n)$-th harmonics that can attain significant amplitudes will be propagative. Thus, this property of the temporal parameters turns them into the best option for adjusting the beamsteering of the system. Therefore, the best choice for setting the reflection and transmission coefficients of the system is selecting the proper configuration (A-G or C-G), as expected. This feature is very interesting for future wireless communications where the communication at both sides of the device is desired [72].
Finally, the efficiency of the proposed analytical method is noteworthy. The computational times associated with our analytical (circuit-based) approach are much lower compared to other numerical techniques. Computation time is less than a second here, while FDTD and other numerical approaches normally take minutes to provide the solution. From a practical perspective, our approach is an efficient method that can be used as a surrogate model to be combined with external optimization or artificial intelligence techniques for an efficient design of future metal-based space-time metastructures.
Although the experimental part is beyond the scope of the paper, it is worth mentioning some of the challenges that a future implementation of the space-time metasurface will face. Naturally, an interesting alternative in the microwaves is the use of PIN and varactor diodes. Some alternatives of this kind can be found in the literature. Switching times and related transients between ON and OFF states in PIN diodes and varactors are indeed one of the main technological challenges. From the theoretical side, we are extending the analytical framework to include the complete circuit model of these active elements [73] (including parasitic effects at higher frequencies). In the low-THz regime, a more exotic alternative could be the use of graphene. Blue and red strips in Fig. 1 could be treated as biased elements at different voltages, thus giving the air and conductor responses. Optical pumping could be an alternative too [74, 75]. The performance of a future prototype will then be validated experimentally in an anechoic chamber and an optic table. The main task will be to check the beamsteering and frequency mixing capabilities of the space-time metasurface by measuring the transmission and reflection coefficients and the diffracted angles with any of the combinations of the three states: air, conductor and grating.
@@ Acknowledgments
This work has been supported by grant TED2021-129938B-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. It has also been supported by grants PID2020-112545RB-C54, PDC2022-133900-I00, and PDC2023-145862-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. It is also part of the grant IJC2020-043599-I funded by MICIU/AEI/10.13039/501100011033 and by European Union NextGenerationEU/PRTR..
Appendix A
The analytical results have been compared with a self-implemented 2D finite-difference time-domain (FDTD) method based on a staggered grid scheme. Two independent codes have been programmed for TE and TM incidences. For an impinging TE-polarized wave ($E_{x},H_{y},H_{z}$), the space-time-varying boundary conditions in the metallic regions of the metasurface (placed in $z_{pos}$) are enforced as $E_{x}=0$ in both conductor and grating states. For TM incidence ($E_{y},E_{z},H_{x}$), the space-time-varying boundary conditions in the metallic regions are enforced as $E_{y}=0$. Second-order Engquist-Majda relations have been implemented as absorbing boundary conditions at the outer limits of the simulation box. The rest of parameters that are involved in the FDTD simulations are: uniform mesh grid with $\Delta y=\Delta z=\frac{\lambda_{0}}{35}$; time step of $\Delta t=C\cdot\frac{\Delta z}{c}$, with $C=0.5$ being the CFL stability condition parameter. To model the zero-thickness metasurface, a homogeneous slab with a thickness less than $0.1\lambda_{0}$ has been considered.